# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ 6. Amenability
Assume that $\mathcal{G}$ has dynamic asymptotic dimension $d$. For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$. Using local compactness, we expand $K$ slightly and write it as $K^{\prime}$, which is open and precompact. Replacing $K^{\prime}$ with $K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, we may assume that $K^{\prime}=K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, which is also open and precompact. Let $\{U_{0},\cdots,U_{d}\}$ be an open cover of $r(K^{\prime})\cup s(K^{\prime})$ with the properties in Proposition 5.8 for the open precompact set $K^{\prime}$ and error estimate $\varepsilon/d+2$. Replacing each $U_{i}$ with its intersection with $s(K^{\prime})\cup r(K^{\prime})$, we may assume that each $U_{i}$ is contained in $s(K^{\prime})\cup r(K^{\prime})\subseteq K^{\prime}$. Claim: $K^{\prime}=\bigcup\limits_{i}K_{i}$, where $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$, $K_{i}$ is open, symmetric, precompact and has compact closure in $\mathcal{G}_{i}$. By Lemma 5.5 (ii) (iii), $s(K^{\prime}\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(1)}\subseteq U_{i}$, thus $K^{\prime}\cap r^{-1}(U_{i}^{(0)})\subseteq K_{i}$. By Lemma 5.5 (i), $\{U_{0}^{(0)},\cdots,U_{d}^{(0)}\}$ covers $\mathcal{G}^{(0)}$. Therefore,
$\bigcup\limits_{i}K_{i}\supseteq\bigcup\limits_{i}K^{\prime}\cap r^{-1}(U_{i}^{(0)})$ $=K^{\prime}\cap r^{-1}(\bigcup\limits_{i}U_{i}^{(0)})$
$=K^{\prime}\cap r^{-1}(\mathcal{G}^{(0)})\supseteq K^{\prime}.$
Clearly, $K_{i}$ is open, precompact and has compact closure in $\mathcal{G}_{i}$, it suffices to show $K_{i}=K_{i}^{-1}$. Indeed, for any $k\in K_{i}$, namely, $k\in K^{\prime}$, $s(g),r(g)\in U_{i}$. Thus $k^{-1}\in K^{\prime}$ as $K^{\prime}=K^{\prime-1}$, and $r(k^{-1})=s(k),s(k^{-1})=r(k)\in U_{i}$, which implies that $k^{-1}\in K_{i}$.
Let $\mathcal{G}_{i}$ be an open subgroupoid of $\mathcal{G}$ generated by the set $K_{i}$, then $\mathcal{G}_{i}$ is amenable follows from Lemma 6.2. Therefore, for each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as in the definition of amenability with respect to the compact subset $\overline{K_{i}}$ and error estimate $\frac{\varepsilon}{d+2}$. Extending $\mu_{i}$ to the whole $\mathcal{G}$ by defining $\mu_{i}(x)=0$ when $x\notin\mathcal{G}_{i}$. We define $\mu:\mathcal{G}\rightarrow[0,1]$, $g\mapsto\sum\limits_{i=0}^{d}\phi_{i}(s(g))\mu_{i}(g)$. Clearly, $\mu$ is a well-defined continuous, compactly supported function in $\mathcal{G}$.
@@
where the third inequality comes from the amenability of $\mathcal{G}_{i}$ and the fact that if $x\in U_{i}$, then $x\in K^{\prime}$ and $s(x)=r(x)=x\in U_{i}$, thus $x\in K_{i}\subseteq\mathcal{G}_{i}$, and if $x\notin U_{i}$, then $\phi_{i}(x)=0$ as supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. The last equality is due to Proposition 5.8.
@@
where the fourth inequality follows from the amenability of $\mathcal{G}_{i}$ and the fact that if $r(k)\in U_{i}$, then $r(k)\in K_{i}$ and $r(r(k))=r(k)$, if $r(k)\notin U_{i}$, then $\phi_{i}(r(k))=0$. The fifth equality is due to the partition of unity.
@@
where the fourth inequality comes from the amenability of $\mathcal{G}_{i}$ and Proposition 5.8. ∎
The following result is inspired by [[Oza12], Theorem 1], which demonstrated that subexponential asymptotic dimension growth for a metric space implies property A. Clearly, slow asymptotic dimension growth also implies property A as it is a subexponential growth. Since asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for the corresponding coarse groupoid (Theorem 4.12), and as property A for a discrete metric space of bounded geometry is equivalent to amenability of the corresponding coarse groupoid [[STY02], Theorem 5.3], we obtain a generalization of the result of Corollary 4.15 in the case of slow growth.
@@
For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$, there exists $R>0$ such that $K\subset K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<R\}$ as the length function $\ell$ is controlled. Note that $K^{\prime}$ is open and precompact, $K^{\prime}=K^{\prime-1}$, $s(K^{\prime})\cup r(K^{\prime})=\mathcal{G}^{(0)}\subseteq K^{\prime}$. By Proposition 5.1, there exists a constant $c$ depending on parameters $\alpha,R,\varepsilon$, a positive integer $p$, and an open cover $\{U_{0},\cdots,U_{f([(R+1)^{c}]R+1)}\}$ of $\mathcal{G}^{(0)}$ with the three properties. Let $\mathcal{G}_{i}$ be an open precompact subgroupoid generated by the set $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$ and error estimate $\frac{\varepsilon}{2}$. Then by Lemma 6.2, $\mathcal{G}_{i}$ is amenable.
For each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as in the definition of amenability with respect to the compact subset $\overline{K_{i}}$ and error estimate $\frac{\varepsilon}{2}$. Extending $\mu_{i}$ to the whole $\mathcal{G}$ by defining $\mu_{i}(x)=0$, if $x\notin\mathcal{G}_{i}$. We define $\mu:\mathcal{G}\rightarrow[0,1],g\mapsto\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(s(g))\mu_{i}(g)$, which is well-defined continuous, compactly supported function on $\mathcal{G}$.
@@
where the third inequality comes from the amenability of $\mathcal{G}_{i}$, and the fourth one is due to the partition of unity.
@@
where the fourth inequality comes from the amenability of $\mathcal{G}_{i}$, and the fifth equality is due to the partition of unity.
@@
where the third equality comes from the equality $\left|a^{p}-b^{p}\right|=\left|a-b\right|\cdot\left|\sum\limits_{j=0}^{p-1}a^{j}b^{p-1-j}\right|$ for all $a,b\geq 0$, the fourth inequality is due to the amenability of $\mathcal{G}_{i}$, and the final inequality follows from Proposition 5.1(3). ∎
@@
Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a strictly increasing function such that $f\preceq x^{\alpha}(0<\alpha<1)$, and let $X=\bigoplus\limits_{n=1}^{\infty}f(n)\mathbb{Z}$, endowed with the direct sum metric. Then $X$ is a discrete metric space with asymptotic dimension growth $f$, a subexponential growth. Note that $X$ has bounded geometry since the sequence $\{f(n)\}$ is (strictly) increasing and every element of $X$ has only finitely many nonzero terms, thus $X$ has property $A$. Hence, the corresponding coarse groupoid $G(X)$ is amenable by Corollary 4.15.
@@ Acknowledgments
We thank Jiawen Zhang for many valuable suggestions on asymptotic dimension growth and Jianchao Wu for useful discussions on coarse groupoids.

### v2

@@ 6. Amenability
Let us assume that $\mathcal{G}$ has dynamic asymptotic dimension $d$. For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$. Using local compactness, we expand $K$ slightly and write it as $K^{\prime}$, which is open and precompact. Replacing $K^{\prime}$ with $K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, we may assume that $K^{\prime}=K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, which remains open and precompact. Next, let $\{U_{0},\cdots,U_{d}\}$ be an open cover of $r(K^{\prime})\cup s(K^{\prime})$, satisfying the properties outlined in Proposition 5.8 for the open precompact subset $K^{\prime}$ and the error estimate $\frac{\varepsilon}{d+2}$. By taking the intersection of each $U_{i}$ with $s(K^{\prime})\cup r(K^{\prime})$, we may assume that each $U_{i}$ is contained in $s(K^{\prime})\cup r(K^{\prime})\subseteq K^{\prime}$. Let $\mathcal{G}_{i}$ represent the subgroupoid generated by the set $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$. According to Proposition 5.8, $\mathcal{G}_{i}$ is both open and precompact.
We claim that $K^{\prime}=\bigcup\limits_{i}K_{i}$, where each $K_{i}$ is open, symmetric, precompact, and has compact closure in $\mathcal{G}_{i}$. We first demonstrate that $K^{\prime}=\bigcup\limits_{i}K_{i}$. According to items (ii) and (iii) of [[GWY17], Lemma 7.4], we have $s(K^{\prime}\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(1)}\subseteq U_{i}$. Thus, it follows that $K^{\prime}\cap r^{-1}(U_{i}^{(0)})\subseteq K_{i}$. Furthermore, by the item (i) of [[GWY17], Lemma 7.4], the collection $\{U_{0}^{(0)},\cdots,U_{d}^{(0)}\}$ covers $\overline{r(K^{\prime})\cup s(K^{\prime})}$. Therefore,
$\bigcup\limits_{i=0}^{d}K_{i}\supseteq\bigcup\limits_{i=0}^{d}K^{\prime}\cap r^{-1}(U_{i}^{(0)})$ $=K^{\prime}\cap r^{-1}(\bigcup\limits_{i=0}^{d}U_{i}^{(0)})$
$\supseteq K^{\prime}\cap r^{-1}(\overline{r(K^{\prime})\cup s(K^{\prime})})\supseteq K^{\prime}.$
Clearly, $K_{i}$ is open, precompact, and has compact closure in $\mathcal{G}_{i}$. It remains to show that $K_{i}=K_{i}^{-1}$. Indeed, for any $k\in K_{i}$, specifically, $k\in K^{\prime}$, $s(g),r(g)\in U_{i}$. Consequently, $k^{-1}\in K^{\prime}$ since $K^{\prime}=K^{\prime-1}$. Furthermore, since $r(k^{-1})=s(k)$ and $s(k^{-1})=r(k)$, we have $r(k^{-1}),s(k^{-1})\in U_{i}$, which implies that $k^{-1}\in K_{i}$. This completes the proof of the claim.
Since $\mathcal{G}_{i}$ is an open precompact subgroupoid of $\mathcal{G}$, it follows from Lemma 6.2 that $\mathcal{G}_{i}$ is amenable. For each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as defined in the context of amenability concerning the compact subset $\overline{K_{i}}$ and the error estimate $\frac{\varepsilon}{d+2}$. We extend $\mu_{i}$ to the entire $\mathcal{G}$ by defining $\mu_{i}(x)=0$ when $x\notin\mathcal{G}_{i}$. We then define a function $\mu:\mathcal{G}\rightarrow[0,1]$, $g\mapsto\sum\limits_{i=0}^{d}\phi_{i}(s(g))\mu_{i}(g)$. Clearly, $\mu$ is a well-defined function that is continuous and has compact support in $\mathcal{G}$.
@@
where the third inequality comes from the amenability of $\mathcal{G}_{i}$ and the fact that if $x\in U_{i}$, then $x\in K^{\prime}$ and $s(x)=r(x)=x\in U_{i}$, thus $x\in K_{i}\subseteq\mathcal{G}_{i}$, and if $x\notin U_{i}$, then $\phi_{i}(x)=0$ as supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. The final equality is due to the item (3) of Proposition 5.8.
@@
where the fourth inequality follows from the amenability of $\mathcal{G}_{i}$ and the fact that if $r(k)\in U_{i}$, then $r(k)\in K_{i}$ and $r(r(k))=r(k)$, if $r(k)\notin U_{i}$, then $\phi_{i}(r(k))=0$. The fifth equality is due to the partition of unity as stated in Proposition 5.8.
@@
where the fourth inequality follows from the amenability of $\mathcal{G}_{i}$ and Proposition 5.8. ∎
The following main theorem is inspired by Corollary 4.19, which shows that for a discrete metric space of bounded geometry $X$, the associated coarse groupoid $G(X)$ with subexponential dynamic asymptotic dimension growth is amenable. It follows that slow dynamic asymptotic dimension growth of the groupoid $G(X)$ guarantees its amenability as this condition indicates subexponential growth. By adopting an approach analogous to that of Theorem 6.3 and using the key technical Proposition 5.1, we can generalize the result of Corollary 4.19 for the case of slow growth.
@@
Without loss of generality, we may assume that $\mathcal{G}$ has dynamic asymptotic dimension function $f\preceq x^{\alpha}$ for $\alpha\in(0,1)$. For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$, there exists $R>0$ such that $K\subseteq K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<R\}$ as the length function $\ell$ is controlled. Note that $K^{\prime}$ is open and precompact, $K^{\prime}=K^{\prime-1}$, and $s(K^{\prime})\cup r(K^{\prime})=\mathcal{G}^{(0)}\subseteq K^{\prime}$. By Proposition 5.1, there exists a constant $c$ depending on parameters $\alpha,R,\varepsilon$, a positive integer $p$ depending only on $\alpha$, and an open cover $\{U_{0},\cdots,U_{f([(R+1)^{c}]R+1)}\}$ of $\mathcal{G}^{(0)}$ with the properties for the open precompact subset $K^{\prime}$ and the error estimate $\frac{\varepsilon}{2}$.
Let $\mathcal{G}_{i}$ be the subgroupoid generated by the subset $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$. Similar to the claim in the proof of Theorem 6.3, we also have the following statement: $K^{\prime}=\bigcup\limits_{i}K_{i}$ and $K_{i}$ is open, symmetric, precompact, and has compact closure in $\mathcal{G}_{i}$. According to Proposition 5.1, $\mathcal{G}_{i}$ is an open precompact subgroupoid of $\mathcal{G}$. Therefore, by Lemma 6.2, $\mathcal{G}_{i}$ is amenable. For each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as in the definition of amenability concerning the compact subset $\overline{K_{i}}$ and the error estimate $\frac{\varepsilon}{2}$. We extend $\mu_{i}$ to the whole $\mathcal{G}$ by defining $\mu_{i}(x)=0$, if $x\notin\mathcal{G}_{i}$. We then define $\mu:\mathcal{G}\rightarrow[0,1],g\mapsto\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(s(g))\mu_{i}(g)$, which is a well-defined continuous, compactly supported function on $\mathcal{G}$.
@@
where the third inequality comes from the amenability of $\mathcal{G}_{i}$, and the fourth one is due to the partition of unity in Proposition 5.1.
@@
where the fourth inequality follows from the amenability of $\mathcal{G}_{i}$, and the fifth equality comes from the partition of unity in Proposition 5.1.
@@
where the third equality comes from the equality $\left|a^{p}-b^{p}\right|=\left|a-b\right|\cdot\left|\sum\limits_{j=0}^{p-1}a^{j}b^{p-1-j}\right|$ for all $a,b\geq 0$, the fourth inequality is due to the amenability of $\mathcal{G}_{i}$, and the final inequality follows from the item (3) of Proposition 5.1. ∎
@@
Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a strictly increasing function such that $f\preceq x^{\alpha}(0<\alpha<1)$, and let $X=\bigoplus\limits_{n=1}^{\infty}f(n)\mathbb{Z}$, which is endowed with the direct sum metric. Then $X$ is a discrete metric space with slow asymptotic dimension growth $f$, a subexponential growth. Moreover, $X$ has bounded geometry since the sequence $\{f(n)\}$ is (strictly) increasing and every element of $X$ has only finitely many nonzero terms. According to Theorem 4.18, $X$ has property $A$. Hence, by Corollary 4.19, the corresponding coarse groupoid $G(X)$ is amenable.
@@ Acknowledgments
We thank Jiawen Zhang for many valuable suggestions on asymptotic dimension growth, Kang Li and Zehong Huang for their helpful comments on dynamic asymptotic dimension, and Jianchao Wu for discussions on coarse groupoids.